# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v2

@@ Abstract
Do home prices incorporate flood risk in the immediate aftermath of specific flood events, or is it the repeated exposure over the years that plays a more significant role? We address this question through the first systematic study of the Italian housing market, which is an ideal case study because it is highly exposed to floods, though unevenly distributed across the national territory. Using a novel dataset containing about 550,000 mortgage-financed transactions between 2016 and 2024, as well as hedonic regressions and a difference-in-difference design, we find that: (i) specific floods do not significantly decrease home prices in areas at risk; (ii) the repeated exposure to floods in flood-prone areas leads to a price reduction, up to 4% in the most frequently flooded regions; (iii) young, low-income buyers are more likely to settle in flood-prone areas, but do not obtain a discount, while more experienced buyers do.
22footnotetext: We would like to thank Paolo Bajardi, Diana Bruni, Matteo Coronese, Paola Danelutti, Adriano di Fabio, Barbara Lenti, Jacopo Lenti, Laura Li Puma, Michela Manuelli, Carla Monferrato, Yamir Moreno, Silvia Ronchiadin, Simone Scarsi, and Fabio Verachi for their helpful comments at various stages of this work, as well as seminar participants at EAERE and IAERE, where this work was selected among the top 5 presentations by young researchers, to be included in a special session at the Italian Economic Society 66th Annual Conference.11footnotetext: Equal contribution.33footnotetext: Corresponding author: marco.pangallo@centai.eu.
@@ 1 Introduction
In this paper, we consider an alternative yet complementary hypothesis, which we term historical memory. We suggest that, in places with storms not as destructive as hurricanes and different housing market institutions than in the U.S., isolated flood events may be insufficient to significantly shift market beliefs and prompt the pricing of flood risk properties. Instead, repeated flood exposure in the same region may have a more substantial influence. If this hypothesis is true, home prices in frequently flooded areas are likely to reflect flood risk more prominently than those in rarely flooded areas, even for properties that remain undamaged. Moreover, experienced buyers, having witnessed more floods, may require a higher discount to purchase at-risk properties compared to younger buyers.
@@
The main reason why there is so little research on how flood risk impacts the Italian housing market is that systematic data only started to exist in the last few years. [^1] We are only aware of one study assessing the impact of flood risk on prices (Rosato et al., 2017), using 197 data points manually collected from real estate agencies in Venice. Only recently systematic data about the housing market became available from the housing advertisement website Immobiliare.it (Pangallo and Loberto, 2018; Loberto et al., 2022), but concerning our research question, these data have only been used to study the impact of a sea wall in Venice (Benetton et al., 2023). In this paper, we take advantage of a novel database that contains the near-universe of mortgages issued by Intesa Sanpaolo, the largest Italian bank. Our final dataset contains about 550,000 mortgages issued from 2016 to 2024, which is about 15% of Italian mortgage-financed transactions, and 10% of total transactions. [^2] The data are highly representative of the Italian housing market throughout the entire national territory, see Supplementary Section A.2. For each transaction, we know the sale price, the date, the income of the buyer, a proxy for whether the buyer is under-36, and several physical characteristics of the transacted home, including its geographical coordinates. We use the coordinates to determine whether a home is inside a flood-prone area by performing spatial matching with risk maps provided by ISPRA, the Italian Institute for Environmental Protection and Research (Trigila et al., 2021). Using the same approach, we also assess whether homes have been hit by specific flood events, such as the 2023 Emilia-Romagna flood. By “hit”, we mean that their coordinates lie within the maximal flooding area identified by the Copernicus EU program (Directorate Space, Security and Migration, European Commission Joint Research Centre , EC JRC).
@@
We find that the 2023 Emilia-Romagna flood (which, as discussed, has been the largest flood in Italy in the last 20 years) did not lead to any further price reduction for homes at risk that were not hit. [^4] We also find that prices of hit homes decreased by 10% in the first 9 months after the event, reversing to a 10% appreciation between 9 and 15 months after the event. Most likely, affected homes were first damaged and then renovated, resulting in a price premium after renovations ended. This result, which contrasts with much of the literature, is robust across specifications, home types (e.g., focusing on ground floor apartments), and considering the second and third largest flood events in the studied period.
@@
We find that homes at risk sell at approximately 1% less than homes that are not at risk. This risk penalty is not uniform. Dividing all transactions in our dataset by our measure of flood risk awareness in the area and at the time of the transaction, we find that flood risk is capitalized in home prices only in the top awareness tercile, where discounts reach 1.7%. The regions which averaged highest awareness over the considered period show stronger discounts, reaching 2.5% in Liguria, 3.5% in Emilia-Romagna and 4% in Toscana.
Third, we study how these results vary across buyers with different age and income. We find that young buyers [^6] More specifically, under-36 buyers with an adjusted income (“ISEE”) less than 40,000€  receiving a mutuo giovani, namely a mortgage with favorable conditions guaranteed by the government. do not obtain any discount for purchasing homes at risk, while more experienced buyers obtain a discount that increases with our awareness measure. This result is an average across all income levels. Further disaggregating by income, we find that income plays a significant role only among young buyers, as high-income young buyers purchase homes at risk with a 3% discount compared to homes that are not at risk, while medium and low-income young buyers do not obtain any discount or even buy at a premium. Finally, in line with Zivin et al. (2023), we find that lower-income buyers are more likely to purchase properties in risky areas, in part because of lower prices.
Summing up, in the immediate aftermath of the largest Italian floods in the last 20 years, we do not observe any price reductions for homes that were in the affected municipalities even though not directly hit. At the same time, we find a small “baseline” penalty of flood risk on prices, but this penalty increases up to four times in the most flood-prone parts of Italy. We interpret this finding as evidence for our hypothesis of “historical memory”, whereby it is not isolated floods, but the repeated exposure to floods, that drives the incorporation of flood risk in home prices. This evidence is further corroborated by our finding that only experienced buyers, who may have witnessed multiple floods in their lifetime, obtain a discount for purchasing homes at risk.
@@
The third piece of literature that we relate to is the one looking at the socio-economic characteristics of the buyers. Although there is substantial evidence that “sophisticated” buyers, such as non-owner-occupiers (Bernstein et al., 2019), or “believers” in climate change (Baldauf et al., 2020), obtain higher discounts when purchasing at-risk properties, the evidence on the effect of age or income is more scarce. For instance, Bradt et al. (2021) find that younger and lower-income buyers are less likely to purchase flood insurance, but Fairweather et al. (2024) do not find any significant difference in terms of searching at-risk properties across income levels. We contribute to this literature by showing interactions between the age and income of the buyers. Regarding the income of buyers who settle in at-risk areas, our main comparison is Zivin et al. (2023), who match information on buyers’ income with transactions. They show that periods in which prices increase correspond to periods in which higher-income buyers are more likely to purchase properties in affected areas. We find a similar result concerning incomes, as we find that areas at risk show slightly lower home prices, thus households with slightly lower income move there. Ellen and Meltzer (2024) also find that lower-income households moved into the areas of New York that were hit harder by Hurricane Sandy.
@@ 2.2 ISPRA risk maps
In this work, we used the 2020 release of the ISPRA maps, which is the most recent update, as a reference for flood risk by the Italian authorities. The most recent tools for hydrological modeling and the inclusion of areas not considered in previous releases were adopted in the 2020 release. Despite these improvements, there are still some limitations in the definition of risk areas due to the uncoordinated work of the basin authorities. The main limitation is the inconsistency between the level of risk. Although in principle ISPRA distinguishes between high, medium, and low risk, classified according to return time, [^18] High-risk areas are subjected to frequent flood events, with return time (i.e., the inverse of the average frequency of occurrence) between 20 and 50 years; medium-risk areas present a return time from 100 to 200 years; low-risk locations are characterized by a return time larger than 200 years (Trigila et al., 2021) this classification is inconsistent across regions. [^19] For example, the high-risk levels are not available for the Marche region (Trigila et al., 2021) in the 2020 release and the notable extension of the high-risk areas in Calabria was inherited from the use of a flood hazard buffer zone across the hydrographic reticulum (Trigila et al., 2021).
Therefore, we build on the ISPRA risk maps to assign a flood risk indicator to each home in the ISP dataset. To accomplish this task, we perform a spatial matching approach, using the coordinates of each housing unit to match it to risk zones. Throughout this work, we introduce the label “at risk” when a home is inside a risk area, without distinction across high, medium, and low levels. While we focus on being at risk, we also perform supplementary analyses on the disaggregated risk levels. From the summary statistics shown in Table LABEL:tab:summary_statistics_categorical, we note that approximately 77% of homes are at no risk of flooding, 11% have low risk, 8% have medium risk, and 4% are at high risk (thus, 23% of homes are in a flood-risk area).
@@ 2.3 Copernicus flooded areas
The Copernicus Emergency Management Service provides geospatial information to support the management of natural disasters, like landslides, wildfires or floods (Directorate Space, Security and Migration, European Commission Joint Research Centre , EC JRC). For flooding events, this data includes detailed vector maps that outline the extent of the flooded area at different times during the emergency.
@@ 2.4 Flood risk awareness: frequency of floods across Italy
To model buyers’ awareness of flood risk we combined historical data of flood events in Italian regions with the existing theoretical insights on flood awareness perception.
@@
Historical data on flood events come from EM-DAT, which is a disaster database that contains all the floods that have occurred in Italy between 2000 and 2023212121Though EM-DAT provides older event records, it discourages the use of the pre-2000 data because of reliability concerns. with at least one among: 10 fatalities; 100 affected people; a declaration of state of emergency; a call for international assistance. These objective inclusion criteria and the popularity of the database make it an ideal starting point for our analysis. As a robustness test, we also considered an alternative dataset that we compiled from ISPRA data (Supplementary Section B.1 provides more details). In both cases, geographical information on floods is provided at the region level, making regions our natural spatial unit for computing awareness over time.
@@
From the theoretical literature on flood risk awareness it is known that perception of risk tends to rise after a new catastrophic event but then it slowly decays as the event is forgotten (Pryce et al., 2011; Tobin and Newton, 1986). Some studies also measure the time scale of awareness. For instance, Heimerl and Meyer (2014) and Egli (2002) report that awareness halves after 7 to 10 years, while Barendrecht et al. (2019) find a 17 years half-life (i.e., the time it takes for a quantity to halve).
@@
We integrate these insights and the actual records of past flood events from EM-DAT to create our estimate of the flood awareness for each Italian region (we cannot go to a lower level of granularity because information on provinces is often lacking or inconsistent). We define the awareness to flood risk in region $r$ at date $t$ as
$a(r,t)=\sum_{d=01/01/2000}^{t}e^{-\frac{(t-d)\cdot\ln{2}}{365\cdot\tau}}\cdot f(r,d),$ (1)
where $f(r,d)$ is 1 if there was a flood in region $r$ at day $d$ according to EM-DAT and 0 otherwise, [^22] If the flood event spans multiple days, we consider the initial day as $d$. We did not consider event severity in our analysis. Although EM-DAT reports multiple severity measures for most events (e.g. fatalities, missing and displaced people, total damage, insured damage, etc), it aggregates this information by events that typically span multiple regions, and this makes it difficult to apportion damages to each region. and $\tau$ is the half-life parameter that represents the time required to halve the influence of each flood event in $a(r,t)$. Thus, the awareness of an Italian region at time $t$ is a weighted sum of the events it experienced, but the contribution of each event decays exponentially as the gap between $t$ and the event time $d$ gets larger. In this way, recent floods contribute more, but their influence decays over time as the event is forgotten.
@@
Overall, $a(r,t)$ combines in a unique variable the number of floods a region has experienced (spatial information) with the time passed between the past floods and the transaction date (temporal information) giving us a good estimate of buyers’ flood risk awareness.
@@
Figure 1 shows the estimated awareness trends per region for the period covered by the ISP data, January 1st, 2016 to August 2024. As the reader can see, starting from 2016, regions show different awareness levels and trends due to the different exposure to floods between 2000 and 2016. For example, we can compare the Emilia-Romagna and Puglia regions. Emilia-Romagna shows a 2.5 awareness level in 2016 and experienced several floods up to August 2024 (spikes in 2018, middle 2019, 2021, and middle 2023) that explain the positive awareness trend. In contrast, Puglia was hit by only one flood in 2003, leading to a low awareness value that decreases in time due to the exponential decay of the weight (Equation 1). [^23] In the main text, we show results considering $\tau=10$ years. 10 years is a value in line with a previous study on the perception of flood risk (Barendrecht et al., 2019). Table A7 in reports the results considering $\tau=7$ and 17 years (coherently with (Barendrecht et al., 2019)), highlighting that our results are robust across different value of $\tau$. A map showing the average awareness of the transaction within each region can be found in Figure 4.
@@
Building on Equation (1), we associate with each transaction in the ISP dataset its awareness at the time of the sale $t_{s}$. In other words, $a(r_{s},t_{s})$ quantifies the level of exposure to past floods that the buyer of each home had experienced when they decided to buy it.

### v3

@@ Abstract
Do home prices incorporate flood risk in the immediate aftermath of specific flood events, or is it the repeated exposure over the years that plays a more significant role? We address this question through the first systematic study of the Italian housing market, which is an ideal case study because it is highly exposed to floods, though unevenly distributed across the national territory. Using a novel dataset containing about 550,000 mortgage-financed transactions between 2016 and 2024, as well as hedonic regressions and a difference-in-difference design, we find that: (i) specific floods do not significantly decrease home prices in areas at risk; (ii) the repeated exposure to floods in flood-prone areas leads to price reductions, up to 4% in the most frequently flooded regions; (iii) low-income buyers are more likely to settle in flood-prone areas.
22footnotetext: We would like to thank Paolo Bajardi, Diana Bruni, Matteo Coronese, Paola Danelutti, Adriano di Fabio, Barbara Lenti, Jacopo Lenti, Laura Li Puma, Michela Manuelli, Carla Monferrato, Yamir Moreno, Silvia Ronchiadin, Simone Scarsi, and Fabio Verachi for their helpful comments at various stages of this work, as well as seminar participants at EAERE and IAERE, where this work was selected among the top 5 presentations by young researchers, to be included in a special session at the Italian Economic Society 66th Annual Conference.11footnotetext: Equal contribution.33footnotetext: Corresponding author: marco.pangallo@gmail.com
@@ 1 Introduction
In this paper, we consider an alternative yet complementary hypothesis, which we term historical memory. We suggest that, in places with storms not as destructive as hurricanes and different housing market institutions than in the U.S., isolated flood events may be insufficient to significantly shift market beliefs and prompt the pricing of flood risk properties. Instead, repeated flood exposure in the same region may have a more substantial influence. If this hypothesis is true, home prices in frequently flooded areas are likely to reflect flood risk more prominently than those in rarely flooded areas, even for properties that remain undamaged.
@@
The main reason why there is so little research on how flood risk impacts the Italian housing market is that systematic data only started to exist in the last few years. [^1] We are only aware of one study assessing the impact of flood risk on prices (Rosato et al., 2017), using 197 data points manually collected from real estate agencies in Venice. Only recently systematic data about the housing market became available from the housing advertisement website Immobiliare.it (Pangallo and Loberto, 2018; Loberto et al., 2022), but concerning our research question, these data have only been used to study the impact of a sea wall in Venice (Benetton et al., 2023) or floods in Emilia-Romagna (Gentili, 2025) (the study by Gentili (2025) is posterior to our work, which was first posted online in February 2025). In this paper, we take advantage of a novel database that contains the near-universe of mortgages issued by Intesa Sanpaolo, the largest Italian bank. Our final dataset contains about 550,000 mortgages issued from 2016 to 2024, which is about 15% of Italian mortgage-financed transactions, and 10% of total transactions. [^2] The data are highly representative of the Italian housing market throughout the entire national territory, see Supplementary Section A.2. For each transaction, we know the sale price, the date, the income of the buyer, a proxy for whether the buyer is under-36, and several physical characteristics of the transacted home, including its geographical coordinates. We use the coordinates to determine whether a home is inside a flood-prone area by performing spatial matching with risk maps provided by ISPRA, the Italian Institute for Environmental Protection and Research (Trigila et al., 2021). Using the same approach, we also assess whether homes have been hit by specific flood events, such as the 2023 Emilia-Romagna flood. By “hit”, we mean that their coordinates lie within the maximal flooding area identified by the Copernicus EU program (Directorate Space, Security and Migration, European Commission Joint Research Centre (EC JRC)., Accessed February 3, 2024).
@@
We find that the 2023 Emilia-Romagna flood (which, as discussed, has been the largest flood in Italy in the last 20 years) did not lead to any further price reduction for homes at risk that were not hit. [^4] We also find that prices of hit homes decreased by 10% in the first 9 months after the event, reversing to a 10% appreciation between 9 and 15 months after the event. Most likely, affected homes were first damaged and then renovated, resulting in a price premium after renovations ended. This result, which contrasts with much of the literature, is robust across specifications, home types (e.g., focusing on ground floor apartments), and considering the second and third largest flood events in the studied period. [^5] This result is also independently confirmed by Gentili (2025), who used a different dataset based on the universe of postings on Immobiliare.it (Loberto et al., 2022).
@@
We find that homes at risk sell at approximately 1% less than homes that are not at risk. This risk penalty is not uniform. Dividing all transactions in our dataset by our measure of flood risk awareness in the area and at the time of the transaction, we find that flood risk is capitalized in home prices only in the top awareness tercile, where discounts reach 1.7%. The regions which averaged highest awareness over the considered period show stronger discounts, reaching 2.5% in Liguria, 3.5% in Emilia-Romagna and 4% in Toscana. Disaggregating discounts by buyers’ socioeconomic characteristics, we further find that only high-income individuals obtain a price discount when purchasing properties at risk, while no comparable effect is observed for low-income buyers.
Third, we examine whether the income of buyers purchasing homes in flood-risk areas differs from that of buyers purchasing in safer locations. In line with Zivin et al. (2023), we find that lower-income households are more likely to buy properties in risky areas. This effect is driven by experienced buyers, while it is absent for young buyers under 36.
Summing up, in the immediate aftermath of the largest Italian floods in the past 20 years, we do not observe any price reductions for homes located in affected municipalities but not directly hit by the event. At the same time, we find a small baseline penalty associated with flood risk, which increases by up to a factor of four in the most flood-prone parts of Italy. We interpret this evidence as supporting our hypothesis of historical memory, whereby it is not isolated floods, but the repeated exposure to floods, that drives the incorporation of flood risk in home prices. In addition, we document income disparities both in the ability to obtain a discount and in the likelihood of settling in flood-risk areas, possibly reflecting differences in bargaining power, education, and financial constraints.
@@
The third piece of literature that we relate to is the one looking at the socio-economic characteristics of the buyers. Although there is substantial evidence that “sophisticated” buyers, such as non-owner-occupiers (Bernstein et al., 2019), or “believers” in climate change (Baldauf et al., 2020), obtain higher discounts when purchasing at-risk properties, the evidence on the effect of age or income is more scarce. For instance, Bradt et al. (2021) find that younger and lower-income buyers are less likely to purchase flood insurance, but Fairweather et al. (2024) do not find any significant difference in terms of searching at-risk properties across income levels. We contribute to this literature by showing interactions between the age and income of the buyers. Regarding the income of buyers who settle in at-risk areas, our main comparison is Zivin et al. (2023), who match information on buyers’ income with transactions. They show that periods in which prices increase correspond to periods in which higher-income buyers are more likely to purchase properties in affected areas. Our results are closest to Ellen and Meltzer (2024), who find that lower-income households moved into the areas of New York that were hit harder by Hurricane Sandy.
@@ 2.2 ISPRA risk maps
In this work, we used the 2020 release of the ISPRA maps, which is the most recent update, as a reference for flood risk by the Italian authorities. The most recent tools for hydrological modeling and the inclusion of areas not considered in previous releases were adopted in the 2020 release. Despite these improvements, there are still some limitations in the definition of risk areas due to the uncoordinated work of the basin authorities. The main limitation is the inconsistency between the level of risk. Although in principle ISPRA distinguishes between high, medium, and low risk, classified according to return time, this classification is inconsistent across regions. [^18] Italian law defines flood hazard classes based on return times (i.e., the inverse of the average frequency of occurrence): high hazard corresponds to return times between 20 and 50 years; medium hazard to return times between 100 and 200 years; and low hazard to return times above 200 years (Trigila et al., 2021). In practice, however, classifications are not harmonized across river basin authorities, even within the same region: many authorities classify as “medium hazard” any area with a return time between 50 and 100 years (https://www.isprambiente.gov.it/pre_meteo/file/Relazioni_metodologiche_mappe_II_ciclo.zip). Additional inconsistencies arise from missing high-hazard data for the Marche region in the 2020 release (Trigila et al., 2021) and from the unusually large extent of high-hazard areas in Calabria, which results from the use of a buffer zone around the hydrographic network rather than a pure return-time criterion (Trigila et al., 2021). Therefore, we will mostly rely on a generic flood risk indicator that does not distinguish between risk levels. Throughout this work, we introduce the label “at risk” when a home is inside a risk area, without distinction across high, medium, and low levels.
To match housing units to ISPRA maps, we perform a spatial matching approach, using the coordinates of each housing unit to match it to risk zones. From the summary statistics shown in Table LABEL:tab:summary_statistics_categorical, we note that approximately 77% of homes are at no risk of flooding, 11% have low risk, 8% have medium risk, and 4% are at high risk (thus, 23% of homes are in a flood-risk area).
@@ 2.3 Copernicus flooded areas
The Copernicus Emergency Management Service provides geospatial information to support the management of natural disasters, like landslides, wildfires or floods (Directorate Space, Security and Migration, European Commission Joint Research Centre (EC JRC)., Accessed February 3, 2024). For flooding events, this data includes detailed vector maps that outline the extent of the flooded area at different times during the emergency.
@@ 2.4 Flood risk awareness: frequency of floods across Italy
To model buyers’ awareness of flood risk, we combined historical data on flood events in Italian regions with theoretical insights from the literature on risk perception.
@@
Historical data on flood events come from EM-DAT, a disaster database that records all floods occurring in Italy between 2000 and 2023202020Although EM-DAT includes earlier records, it advises against using pre-2000 data due to reliability concerns.. Events are included if they satisfy at least one of the following criteria: 10 fatalities; 100 affected people; a declaration of state of emergency; a call for international assistance. The objectivity of these inclusion criteria, together with the widespread use of the database, makes EM-DAT an ideal starting point for our analysis. As a robustness check, we also consider an alternative dataset that we compiled from ISPRA data (see Supplementary Section B.1 for details). In both datasets, geographic information is provided at the regional level, making regions the natural spatial unit for computing awareness over time.
@@
The theoretical literature on flood-risk awareness shows that risk perception typically rises after a major event and then gradually declines as the event fades from memory (Pryce et al., 2011; Tobin and Newton, 1986). Several studies also estimate the time scale of this decay. For example, Heimerl and Meyer (2014) and Egli (2002) report that awareness halves after 7/10 years respectively, while Barendrecht et al. (2019) find a half-life of 17 years (i.e., the time required for the perceived risk to diminish by half).
@@
We integrate these theoretical insights with the actual record of past flood events from EM-DAT to construct an estimate of flood awareness for each Italian region. (We cannot work at a finer spatial scale because information at the province level is often incomplete or inconsistent.) We define the awareness of flood risk in region $r$ on day $t$ as
$a(r,t)=\sum_{d=t_{0}}^{t}2^{-\frac{(t-d)}{\tau}}\cdot\mathcal{I}(r,d),$ (1)
where the indicator variable $\mathcal{I}(r,d)=1$ if EM-DAT reports a flood in region $r$ on day $d$ (and 0 otherwise), [^21] If a flood spans multiple days, we take the first day as $d$. and $\tau$ is the half-life parameter, representing the number of days required for the influence of a flood event to halve. Thus, the awareness in a region at time $t$ is a weighted sum of the flood events it has experienced, with each event’s contribution decaying exponentially as the time since the event increases. [^22] We set $t_{0}=$ 01/01/2000 for the reasons discussed in Footnote 20. In this way, recent floods matter more, but their influence diminishes steadily as they are forgotten.
@@
Overall, $a(r,t)$ combines in a single variable both the number of floods a region has experienced (spatial information) and the time elapsed between each flood and the sale date (temporal information), yielding a simple measure of buyers’ awareness of flood risk.
Our measure is intentionally as assumption-free as possible, with $\tau$ as the only degree of freedom. All our analyses are robust across the three values $\tau=7,10,17$ years reported in the literature (Heimerl and Meyer, 2014; Egli, 2002; Barendrecht et al., 2019). (We report results for the intermediate value $\tau=10$ in the main text and provide results for the other values in the Appendix.) Another advantage of this formulation is that it captures awareness rather than realized damages: it applies to the entire regional territory (most of which has not experienced floods), it spans a long pre-sample period (2000–2016), and it does not rely on event severity, which is often inconsistently recorded and difficult to match to regions in our data.
@@
As an illustration, Figure 1 shows the estimated awareness trends by region for the period covered by the ISP data (January 1st, 2016 to August 2024). Regions start the sample with different awareness levels due to their varying exposure to floods between 2000 and 2016. For example, Emilia-Romagna shows an awareness level of about 2.5 in 2016 and experienced several floods up to August 2024 (spikes in 2018, mid-2019, 2021, and mid-2023), producing an upward trend. In contrast, Puglia experienced only one flood in 2003, resulting in a low awareness level that steadily declines over time due to the exponential decay defined in Equation 1.
@@
Building on Equation (1), we associate with each transaction in the ISP dataset the awareness level at the time of sale $t_{s}$. That is, $a(r_{s},t_{s})$ quantifies the level of exposure to past floods that the buyer of each home faced at the moment of purchase.